# Unleashing the Potential of Large Language Models for Predictive Tabular Tasks in Data Science

### v2

@@ 1 Introduction
This work seeks to explore the potential of LLMs in comprehending and leveraging the relational and semantic richness of tabular data through large-scale, table-specific pretraining. To accomplish this objective, we have compiled a comprehensive pretraining corpus that includes a vast array of tables from diverse domains. The tables are mainly sourced from Kaggle, that contains massive tabular data in CSV for data science, consisting of around 300 domains. This offers our model with exposure to a wide range of tabular structures and content, thus enhancing its ability to generalize across different types of data, and facilitating the model’s understanding of the relationships within the tabular data. Furthermore, we propose a straightforward approach by adopting the self-supervised learning paradigm, inherent in LLMs, applied to our curated collection of tabular data, aiming to unearth intricate data patterns, automate feature engineering processes, and refine predictive modeling techniques. Specifically, we introduce a unified training framework that seamlessly integrates table contents with task-specific instructions, enabling the execution of various training tasks and fostering reasoning between the provided instructions and the tabular data. This study aims to explore the potential of Large Language Models (LLMs) in comprehending and leveraging the relational and semantic richness inherent in tabular data through large-scale, table-specific pretraining. To achieve this objective, we have compiled an extensive pretraining corpus comprising a diverse array of tables sourced primarily from Kaggle, which hosts vast quantities of tabular data in CSV format across approximately 300 domains. This corpus provides our model with exposure to a wide spectrum of tabular structures and content, thereby enhancing its capacity to generalize across different data types and facilitating a deeper understanding of the relationships within the tabular data.
Additionally, we propose a straightforward approach by employing the self-supervised learning paradigm inherent in LLMs, applied to our curated collection of tabular data. This approach aims to uncover intricate data patterns, automate feature engineering processes, and refine predictive modeling techniques. Specifically, we introduce a unified training framework that seamlessly integrates table contents with task-specific instructions, enabling the execution of various training tasks and fostering reasoning between the provided instructions and the tabular data. With the innate human intent understanding capabilities offered by LLMs, our pretraining approach further equips LLMs with the combined benefits of table comprehension and natural language understanding, enabling them to effectively tackle challenging predictive tasks associated with tabular data.
@@
Our trained model demonstrates exceptional performance, substantiated through extensive experimental analysis and comparative assessments across 30 classification and regression tasks. Compared to the Llama-2 baseline, our approach achieves an average improvement of 8.9% in classification tasks and 10.7% in regression tasks. For missing value prediction tasks, our model outperforms GPT-4 by 27%. Furthermore, our model exhibits a significant 28.8% improvement in extreme-few-shot (4-shot) predictions on diverse datasets and a notable 18.8% progress in tasks involving extensive context learning. In the context of extensive context learning, our model shows a remarkable performance increase of 25.9% over Llama-2 80K (Fu et al., 2024).

### v4

@@ 1 Introduction
This study aims to explore the potential of LLMs in comprehending and leveraging the relational and semantic richness inherent in tabular data through large-scale, table-specific pretraining. To achieve this objective, we have compiled an extensive pretraining corpus comprising a diverse array of tables sourced primarily from Kaggle, which hosts vast quantities of tabular data in CSV format across approximately 300 domains. This corpus provides our model with exposure to a wide spectrum of tabular structures and content, thereby enhancing its capacity to generalize across different data types and facilitating a deeper understanding of the relationships within the tabular data. Furthermore, we propose a straightforward approach by adopting the self-supervised learning paradigm, inherent in LLMs, applied to our curated collection of tabular data, aiming to unearth intricate data patterns, automate feature engineering processes, and refine predictive modeling techniques. Specifically, we introduce a unified training framework that seamlessly integrates table contents with task-specific instructions, enabling the execution of various training tasks and fostering reasoning between the provided instructions and the tabular data. With the innate human intent understanding capabilities offered by LLMs, our pretraining approach further equips LLMs with the combined benefits of table comprehension and natural language understanding, enabling them to effectively tackle challenging predictive tasks.
@@
Our trained model demonstrates impressive performance, substantiated through extensive experimental analysis and comparative assessments across 30 classification and regression tasks. Compared to the Llama-2, our approach achieves an average improvement of 8.9% in classification tasks and 10.7% in regression tasks. For missing value prediction tasks, our model outperforms GPT-4 by 27%. Furthermore, our model exhibits a significant 28.8% improvement in extreme-few-shot (4-shot) predictions on diverse datasets and a notable 18.8% progress in tasks involving extensive context learning. In the context of extensive context learning, our model shows a remarkable performance increase of 25.9% over Llama-2 80K (Fu et al., 2024).